Consider the following hypothesis test: 38. O 39. 40. O Ho: p 2 0.75 Ha:p<0.75 A sample of 400 items was selected. Compute the p-value and state your conclusion for each of the following sample results. Use a- 0s Round your answers to four decimal places. . P-0.68 p-value p-value Select b. p-0.73 p-value Conclusion: p-value Select p-value Conclusion: p-value (Select Ho d, p=0.77 p-value Conclusion: p-value (Select Ho
Answer #1

a)

null Hypothesis alternate Hypothesis: Ho: 0.750 Ha 0.750 for 0.05 level with left tailed test, critical value of z- Decision rule: 1.645 reject Ho if test statisticz < -1.645 sample success 272 sample size 400.0 std error -Se p( 1-p)/n) = 0.0217 sample proportion p 0.6800 test stat = z (p-p)/Se= 3.23 p value 0.0006

conclusion: p value <0.05 ; reject Ho

b)

p value =0.1788 ( try 0.1778 if this comes wrong due to rounding)

conclusion: p value >0.05 ; fail to reject Ho

c)

p value =0.0104 ( try 0.0105 if this comes wrong due to rounding)

conclusion: p value <0.05 ; reject Ho

d)

p value =0.8222 ( try 0.8212 if this comes wrong due to rounding)

conclusion: p value >0.05 ; fail to reject Ho